OpenStudy (anonymous):

Calculate the following integral. $(z+2)(rt(1-z))dz

hartnn (hartnn):

you mean \[\int \dfrac{z+2}{\sqrt{1-z}}dz\]

hartnn (hartnn):

or, \[\int {(z+2)}{\sqrt{1-z}}dz \]

OpenStudy (anonymous):

This is a problem from a quiz concerning the section on u-substitution. I got it wrong, so maybe you guys can help me figure it out! @hartnn - no, I do not, I mean for the second term to be also on top.

OpenStudy (anonymous):

The second one is what I mean.

hartnn (hartnn):

right, so put \(\sqrt{1-z}=u \\ then, z=1-u^2\) dz=..?

OpenStudy (anonymous):

Wouldn't it be simpler to set u=1-z? du=(1/(2((1-z)^(-1/2)))(-1)dz

hartnn (hartnn):

simpler ? no.... du=(1/(2((1-z)^(-1/2)))(-1)dz itself looks complicated compared to dz = -2u du

OpenStudy (anonymous):

Quite true...

hartnn (hartnn):

so, your new integral after substitution will be.... ?

OpenStudy (anonymous):

-2$(3u-u^3)(rt(u))du?

hartnn (hartnn):

how come root sign come again! ? that substitution was done to eliminate the \(\sqrt . \) sign...

OpenStudy (anonymous):

Oh, whoops! No root. But is the rest correct?

hartnn (hartnn):

-2u (3-u^2)u du yes, thats correct :)

hartnn (hartnn):

u know how to proceed ?

OpenStudy (anonymous):

Yep! Thank you!

hartnn (hartnn):

welcome ^_^